# Utilize the Flow before Stepping into the Same River Twice: Certainty Represented Knowledge Flow for Refusal-Aware Instruction Tuning

### v1

@@ Abstract
Refusal-Aware Instruction Tuning (RAIT) enables Large Language Models (LLMs) to refuse to answer unknown questions. By modifying responses of unknown questions in the training data to refusal responses such as “I don’t know”, RAIT enhances the reliability of LLMs and reduces their hallucination. Generally, RAIT modifies training samples based on the correctness of the initial LLM’s response. However, this crude approach can cause LLMs to excessively refuse answering questions they could have correctly answered, the problem we call over-refusal. In this paper, we explore two primary causes of over-refusal: Static conflict emerges when the RAIT data is constructed solely on correctness criteria, causing similar samples in the LLM’s feature space to be assigned different labels (original vs. modified “I don’t know”). Dynamic conflict occurs due to the changes of LLM’s knowledge state during fine-tuning, which transforms previous unknown questions into knowns, while the training data, which is constructed based on the initial LLM, remains unchanged. These conflicts cause the trained LLM to misclassify known questions as unknown, resulting in over-refusal. To address this issue, we introduce Certainty Represented Knowledge Flow for Refusal-Aware Instructions Construction (CRaFT). CRaFT centers on two main contributions: First, we additionally incorporate response certainty to selectively filter and modify data, reducing static conflicts. Second, we implement preliminary rehearsal training to characterize changes in the LLM’s knowledge state, which helps mitigate dynamic conflicts during the fine-tuning process. We conducted extensive experiments on open-ended question answering and multiple-choice question task. Experiment results show that CRaFT can improve LLM’s overall performance during the RAIT process. Source code and training data will be released at Github.
@@ 3.2 Static Conflicts in Cor-RAIT
Figure 4: RAIT Data Construction.
@@ 3.3 Certainty Mitigates the Static Conflicts
Theoretical analysis (see Appendix A.2) indicates a weak (non-differentiable) link between the LLM’s feature and the response correctness $\mu$ for the specific question $q$. This weak correlation causes highly similar samples being categorized into $D_{\text{van}}$ and $D_{\text{idk}}$ respectively. To mitigate this, we propose incorporating a robust indicator variable aligned with correctness to select and construct the RAIT data. This variable should ensure that similar samples share comparable values, reducing the above misclassification. We suggest adopting the certainty of the LLM’s response as the indicator variable. Our theoretical analysis (see Appendix A.3) shows that certainty meets the above requirements. [^3] Section 4.2 discusses various methods for representing LLM response certainty. Appendix A.3 uses entropy as a measure, but our findings extend to other methods as well.
@@ 3.4 Knowledge Flow and Dynamic Conflict
We use rehearsal training (see Chapter 4.3) to capture the LLM’s natural knowledge flow during SFT. The experiment on MMLU dataset (Hendrycks et al. 2020) and LLaMA-3-8B-instruct showed that the correctness of 69% samples, initially below 0.5, improved, thereby validating the above analysis on dynamic conflict. More experimental results are in Appendix A.5.
@@ 4.1 Overview
Figure 6: The Framework of CRaFT.
@@
[…]
@@
Using the knowledge state and flow from Stage 1, we select suitable samples from $D_{\text{src}}$ to construct the RAIT data $D_{\text{RAIT}}$ (see Section 4.4), which is is used to fine-tune the initial model $\mathcal{M}_{0}$.
4.2 Knowledge State Query
@@
4.3 Rehearsal Training and Knowledge Flow

### v2

@@ Abstract
Refusal-Aware Instruction Tuning (RAIT) enables Large Language Models (LLMs) to refuse to answer unknown questions. By modifying responses of unknown questions in the training data to refusal responses such as “I don’t know”, RAIT enhances the reliability of LLMs and reduces their hallucination. Generally, RAIT modifies training samples based on the correctness of the initial LLM’s response. However, this crude approach can cause LLMs to excessively refuse answering questions they could have correctly answered, the problem we call over-refusal. In this paper, we explore two primary causes of over-refusal: Static conflict occurs when similar samples within the LLM’s feature space receive differing supervision signals (original vs. modified “I don’t know”). Dynamic conflict, on the other hand, emerges as the LLM’s knowledge evolves during SFT, allowing it to answer questions that were previously unanswerable. Yet, these now-answerable training samples still retain the original “I don’t know” supervision signals based on the initial LLM state, resulting in inconsistencies. These conflicts cause the trained LLM to misclassify known questions as unknown, resulting in over-refusal. To address this issue, we introduce Certainty Represented Knowledge Flow for Refusal-Aware Instructions Tuning (CRaFT). CRaFT centers on two main contributions: First, we additionally incorporate response certainty to selectively filter and modify data, reducing static conflicts. Second, we implement preliminary rehearsal training to characterize changes in the LLM’s knowledge state, which helps mitigate dynamic conflicts during the fine-tuning process. We conducted extensive experiments on open-ended question answering and multiple-choice question task. Experiment results show that CRaFT can improve LLM’s overall performance during the RAIT process. Source code and training data will be released at Github.
@@ 1 Introduction
In conducting our experimental analysis, we sought a well-founded metric within current research. However, existing methods have notable limitations. For instance, (Feng et al. 2024) proposed using multiple evaluation metrics simultaneously, which is challenging to manage as it is difficult for a model to optimize across multiple criteria at once. Additionally, we provided counterexamples to demonstrate the inadequacy of metrics proposed in other studies. Consequently, we examined these shortcomings and introduced a singular and comprehensive metric: Truthful Helpfulness Score(THS).
@@ 3.2 Static Conflicts in Cor-RAIT
Figure 4: RAIT Data Construction of Cor-RAIT, CorCer-RAIT and CRaFT. Cor-RAIT partitions data based on accuracy $\mu$ and a threshold $\tau_{\mu}$. For CorCer-RAIT, $D_{\text{van}}$ is derived from samples with accuracy exceeding the threshold and the highest certainty, while $D_{\text{idk}}$ consists of samples with accuracy below the threshold and the lowest certainty. CRaFT employs a two-stage process: in the first stage, data with $\Delta\mu>0$ is excluded through knowledge queries; the second stage follows the same procedure as CorCer-RAIT.
@@ 3.3 Certainty Mitigates the Static Conflicts
We conducted a theoretical analysis 333detailed proof in Appendix A.2 establishing a weak (non-differentiable) link between the LLM’s feature and the response correctness $\mu$ for the specific question $q$. This weak correlation causes highly similar samples being categorized into $D_{\text{van}}$ and $D_{\text{idk}}$ respectively. To mitigate this, we propose incorporating a robust indicator variable aligned with correctness to select and construct the RAIT data. This variable should ensure that similar samples share comparable values, reducing the above misclassification. We suggest adopting the certainty of the LLM’s response as the indicator variable. Our theoretical analysis shows that certainty meets the above requirements. [^4] Section 4.2 discusses various methods for representing LLM response certainty. Appendix A.3 uses entropy as a measure, but our findings extend to other methods as well.
@@ 3.4 Knowledge Flow and Dynamic Conflict
We select the data with the highest correctness and certainty for SFT, a process we refer to as rehearsal training [^5] Details are provided in Appendix B.2. Rehearsal training is designed to capture the LLM’s natural knowledge flow during SFT. Experiments on the MMLU dataset (Hendrycks et al. 2020) and LLaMA-3-8B-instruct demonstrated that the correctness of 69% of samples, initially below 0.5, improved, thereby validating the aforementioned analysis on dynamic conflict. Additional experimental results are provided in Appendix A.5.
@@ 4.1 Overview
Figure 6: The Framework of CRaFT: Stage 1 queries the knowledge state and flow, while Stage 2 constructs and tunes refusal-aware instructions.
@@
The output of stage one is the knowledge state and flow indicators of the model. First, we perform a knowledge state query to obtain the correctness and certainty of the model’s responses to the samples in the source dataset. Next, we conduct rehearsal training on the model, resulting in the perturbed version. By comparing the knowledge states before and after perturbation, we derive the indicators of knowledge flow during the supervised fine-tuning process.
@@
Using the knowledge state and flow from Stage 1, we select suitable samples from $D_{\text{src}}$ to construct the RAIT data, which is is used to fine-tune the initial model.
4.2 Query the Knowledge State and Flow of LLM
Knowledge State Query
@@ 4.2 Query the Knowledge State and Flow of LLM
Rehearsal Training and Knowledge Flow